# Feature Attribution with Necessity and Sufficiency via Dual-stage Perturbation Test for Causal Explanation

### v1

@@ B.1 Datasets and Target Models
  • CIFAR10(Krizhevsky et al., 2009) contains colour images of the 10 categories.
@@ B.2 Baselines
  • Saliency(Simonyan et al., 2014) construct contributions using absolute values of partial derivatives.
@@
  • IDGI(Yang et al., 2023b) enhances attribution quality by removing noise in IG-based methods through focused use of gradients along the important direction.
@@ B.3 Evaluation Metrics
  • Infidelity (INF)(Yeh et al., 2019) measures the expected mean square error between the dot product of an attribution and input perturbation, and the difference of output after significant perturbation associated with the attribution.
  • IROF (IR)(Rieger & Hansen, 2020) calculates the area under the curve per class based on the sorted mean importance of iteratively removed feature segments.
  • ${}^{+}$ (FID+) and ${}^{-}$ (FID-)(Amara et al., 2022) respectively evaluate the proportions of predictions changing and remaining unchanged when features with high attribution scores are removed and retained.
  • Max-Sensitivity (MS)(Yeh et al., 2019) computes the maximum difference between new and original attributions, where the new attributions correspond to perturbed inputs.
  • Sparseness (SPA)(Chalasani et al., 2020) applies the Gini index to the absolute values on the attribution.
@@
B.4 Implementation Details
For the parametric explanation methods (MP, GNNExplainer, PGExplainer, ReFine, CFGNExplainer, CIMI, MixUpExplainer), we conducted a grid search to tune their hyperparameters. For our FANS, we utilized the Adam optimizer and set the learning rate to 0.001 for image datasets and 0.1 for graph datasets. The maximum number of training epochs was set to 50 for image datasets and 30 for graph datasets. The times of perturbation $t$ in Eq. 11 and $u$ in Eq. 13 were both set to 50. while the resampling sizes were set to 1.

### v3

@@ C.1 Datasets and Target Models
  • CIFAR10 (Krizhevsky et al., 2009) contains color images of the 10 categories.
@@ C.2 Baselines
  • Saliency (Simonyan et al., 2013) construct contributions using absolute values of partial derivatives.
@@
  • IDGI (Yang et al., 2023) enhances attribution quality by removing noise in IG-based methods through the focused use of gradients along the important direction.
@@
  • CF2 (Tan et al., 2022) proposes an objective function for optimizing the edge mask, which consists of factual loss, counterfactual loss, and $L_{1}$ regularization.
  • NSEG (Cai et al., 2022) uses SCMs for counterfactual probability estimation and optimizes continuous masks via gradient updates, using the PNS lower bound as the objective to generate necessary and sufficient GNN explanations.
@@ C.3 Evaluation Metrics
C.3.1 Infidelity (INF)
The Infidelity (INF) metric (Yeh et al., 2019) quantifies the expected mean square error between the dot product of an attribution vector and the difference of input and perturbed input, and the difference of output before and after perturbations. Mathematically, Infidelity is defined as follows.
$\small\mathrm{INFD}(\mathbf{s},\mathbf{x}^{t},f)=\mathbb{E}_{\mathbf{m}\sim U(\mathbf{0},\mathbf{1})}\left[\left((\mathbf{x}^{t}-\mathbf{x}^{t}\circ\mathbf{m})^{T}\mathbf{s}-\left(f(\mathbf{x}^{t})-f(\mathbf{x}^{t}\circ\mathbf{m})\right)\right)^{2}\right],$
where $U(\mathbf{0},\mathbf{1})$ denotes the uniform distribution over the hypercube $[0,1]^{D}$ in $D$-dimensional space, $\mathbf{s}\in\mathbb{R}^{D}$ represent the attribution vector determined by a FA method.
C.3.2 IROF (IR)
Iterative Removal Of Features (IROF or IR) (Rieger & Hansen, 2020) calculates the Area Over the Curve (AOC) for the class score based on the sorted mean importance of feature segments as they are iteratively removed. Formally, IROF is defined as follows.
$\text{IROF}(\mathbf{a}_{1},...,\mathbf{a}_{N},\mathbf{x}^{t}_{1},...,\mathbf{x}^{t}_{N},f)=\frac{1}{N}\sum_{n=1}^{N}\text{AOC}\left(\frac{f(t(\mathbf{x}^{t}_{n},\mathbf{a}_{n},1))_{i}}{f(\mathbf{x}^{t}_{n})_{i}},...,\frac{f(t(\mathbf{x}^{t}_{n},\mathbf{a}_{n},L))_{i}}{f(\mathbf{x}^{t}_{n})_{i}}\right),$
where $i$ is a class index, $L$ is the number of segments, $\mathbf{x}^{t}_{n}$ represents the $n$-th target input to be explained, $\mathbf{a}_{n}$ denotes the attribution of $\mathbf{x}^{t}_{n}$ predicted by the feature attribution method to be evaluated. $t(\mathbf{x}^{t}_{n},\mathbf{a}_{n},l)$ is designed to perturb the input $\mathbf{x}^{t}_{n}$ in a segment-wise manner. IROF calculates the importance of each segment in $\mathbf{x}^{t}_{n}$ by averaging the significance scores of individual features within those segments, as provided by its attribution $\mathbf{a}_{n}$. The segments are then sorted based on their importance scores in descending order. Function $t$ selectively perturbs the $l$-th ranked segment, yielding the modified input as its output.
C.3.3 Fidelity+ (FID+) and Fidelity- (FID-)
Fidelity+ metric (FID+) and Fidelity- metric (FID-) (Amara et al., 2022) respectively evaluate the proportions of predictions changing and remaining unchanged when features with high attribution scores are removed and retained. For the sake of convenience in discussion, we assume that $f$ is a multi-class classification model.
$\small\operatorname{Fidelity}^{+}(\mathbf{s}_{1},...,\mathbf{s}_{N},\mathbf{x}^{t}_{1},...,\mathbf{x}^{t}_{N},f)=1-\frac{1}{N}\sum_{i=1}^{N}\mathbb{I}\left(\arg\max_{i}f(\mathbf{x}^{t}_{i})=\arg\max_{i}f(\mathbf{x}^{t}_{i}\circ(\mathbf{1}-\mathbf{s}))\right),$ (16)
$\small\operatorname{Fidelity}^{-}(\mathbf{s}_{1},...,\mathbf{s}_{N},\mathbf{x}^{t}_{1},...,\mathbf{x}^{t}_{N},f)=1-\frac{1}{N}\sum_{i=1}^{N}\mathbb{I}\left(\arg\max_{i}f(\mathbf{x}^{t}_{i})=\arg\max_{i}f(\mathbf{x}^{t}_{i}\circ\mathbf{s})\right),$ (17)
where $\arg\max_{i}f(\mathbf{x})$ denotes the category with the maximum predicted probability for a given input vector $\mathbf{x}$, where $f(\mathbf{x})$ represents the model’s output, typically a vector of class probabilities.
C.3.4 Max-Sensitivity (MS)
Max-Sensitivity (MS) (Yeh et al., 2019) computes the maximum difference between new and original attributions, where the new attributions correspond to perturbed inputs. Given an input neighborhood radius $r$ and an attribution method $h$, the max-sensitivity is defined as:
$\operatorname{SENS}_{\text{MAX }}(h,\mathbf{x}^{t},f,r)=\max_{\|\mathbf{z}-\mathbf{x}^{t}\|\leqslant r}\|h(f,\mathbf{z})-h(f,\mathbf{x}^{t}))\|.$ (18)
C.3.5 Sparseness (SPA)
The Sparseness (SPA) metric (Chalasani et al., 2020) is utilized to quantify the sparsity of an attribution, leveraging the Gini index on the absolute values of the attribution scores. Specifically, given a feature attribution $\mathbf{s}^{\text{sort}}$ sorted in descending order, the Gini Index is given by
$\operatorname{Gini}(\mathbf{s}^{\text{sort}})=1-2\sum_{d=1}^{D}\frac{s_{d}}{\|\mathbf{s}\|_{1}}\left(\frac{D-d+0.5}{D}\right)$ (19)
C.3.6 Recall@N
@@
Figure 7: Efficiency analysis on the CIFAR10 dataset.
Figure 8: Parameter sensitivity analysis for FANS in Eq. 11 and Eq. 13.: (a)-(b) boundary $b$ of the neighborhood of the target input $\mathbf{x}^{t}$. (c)-(d) threshold $c$ for prediction change. The triangle $\nabla$’ represents the values of $b$ or $c$ that we obtained using our heuristic strategy.
C.4 Implementation Details
For the explanation methods with hyperparameters (MP, GNNExplainer, PGExplainer, ReFine, CFGNExplainer, CIMI, MixUpExplainer), we conducted a grid search to tune their hyperparameters. We choose the most commonly used baseline settings based on the type of data. Specifically, for the image dataset, we assign a value to $\mathbf{x}^{\prime}$ by sampling from a uniform distribution defined on $[0,1]^{d}$. For the graph dataset, we fix $\mathbf{x}^{\prime}$ to an all-zero vector.
Below are the details of our FANS. Since FANS belong to a feature subset-wise attribution method, we demonstrate the effectiveness of FANS by evaluating the feature subset with the highest PNS estimated by FANS through optimization w.r.t. dimension subset $\mathbf{s}$ (i.e., we adopt the setting of FANS introduced in Section 4.7). In practice, we utilized the Adam optimizer and set the learning rate to 0.001 for image datasets and 0.1 for graph datasets. The maximum number of training epochs was set to 50 for image datasets and 30 for graph datasets. The times of perturbation $t_{\textsc{SF}}$ in Eq. 12 and $t_{\textsc{NC}}$ in Eq. 14 were both set to 50. while the resampling sizes were set to 1. According to our proposed strategy in Section 4.6, the values of boundary $b$ on MNIST, Fashion-MNIST, CIFAR10, BACommunity, Citeseer, and Pubmed are set to $1.0539$, $1.0534$, $1.0546$, $1.0599$, $1.0599$ and $1.0600$, respectively. Correspondingly, the threshold $c$ are set to $0.0534$, $0.0538$, $0.0118$, $7.0710\times 10^{-6}$, $1.3271\times 10^{-6}$ and $4.0706\times 10^{-7}$ for the above datasets, respectively.
Appendix D Efficiency Analysis
We evaluate the efficiency of FANS on CIFAR10, in terms of the inference time. We select representative baselines, including optimization-based methods MP, LIME, and CIMI, as well as non-optimization-based methods FeatAblation, Occlusion, DeepLift, and IG. As shown in Figure 7, the following two main conclusions can be drawn. First, for the same sample size, FANS remains competitive in terms of inference time. Second, for the same time, when the time is approximately greater than 10 seconds, FANS can explain more samples than MP and LIME in the same amount of time. For instance, when given an inference time of 100 seconds, FANS can process approximately 450 samples, whereas LIME and MP can only handle around 350 and 75 samples, respectively. Note that the efficiency of FeatAblation, Occlusion, DeepLift, and IG remains stable, which is expected since they simply compute gradients or perturbations for each sample.
Appendix E Sensitivity Analysis of boundary $b$ and threshold $c$
We present the sensitivity analysis for the parameters $b$ and $c$ in our FANS. Figure 8 (a) and (b) show that the performance of FANS is not sensitive to changes in the values of $b$. In particular, values of $b=1.055$ on CIFAR10 and $b=1.053$ on MNIST computed by our proposed heuristic strategy in Section 4.6 typically give a robust and comparable performance. For the value of $c$, the performance of FANS gradually decreases and becomes stable, with the value of $c=0.012$ on CIFAR10 and $c=0.053$ on MNIST also giving comparable performance.